1. In a game of bowls one particular bowl hits the jack head on causing it to move forward. The jack has a mass of 300 g and was originally stationary.  The bowl has a mass of 1 kg and was moving at a speed of 2 m/s before hitting the jack, but continued forwards at 1.5 m/s after the collision.  What is the speed of the jack after the collision?



  1. Police suspect that a van was breaking the speed limit when it collided with a parked car. Investigations provided them with the following data: mass of parked car   = 800 kg,  mass of van = 1000 kg,  velocity of  car after collision = 5 m/s,  velocity of van after collision = 11 m/s


Show whether the van was breaking the 30 mph speed limit (equivalent to 13 m/s



                           











1
Expert's answer
2021-04-26T16:48:25-0400

1)


"mu=mu'+Mv\\(1)(2-1.5)=0.3v\\v=1.67\frac{m}{s}"

2)


"mu=mu'+Mv\\(1000)(u-11)=800(5)\\u=15\frac{m}{s}"

Yes, the van was breaking the 30 mph speed limit .